### v2

@@ 1 Introduction
In this paper, we present NBDI (Novelty-based Decision Point Identification) [^1] Our code is available at:  https://anonymous.4open.science/r/nbdi-4D44, a simple state-action novelty-based decision point identification method that allows the agent to learn terminated skills from task-agnostic demonstrations. In this context, the term task-agnostic refers to the collection of trajectories from a diverse set of tasks, excluding the one we are specifically interested in (see Appendix E for visualizations). Identifying critical decision points promote knowledge transfer between different tasks and stimulate exploration by closely connecting different areas in the state space (McGovern & Barto, 2001; Menache et al., 2002; Şimşek & Barto, 2004). For example, detecting doorways between rooms is useful regardless of the specific task at hand. We demonstrate the straightforward applicability of our method to the skill-based deep RL framework, and illustrate how it can lead to improvements in decision-making. To summarize, our three main contributions are as follows: (i) For the first time, we propose a skill termination condition for task-agnostic demonstrations that remains effective even when the environment configuration of complex, long-horizon downstream tasks undergo significant changes. (ii) We present a novel finding in reinforcement learning, which is the identification of state-action novelty-based critical decision points. Furthermore, we demonstrate that executing terminated skills, based on state-action novelty, can substantially enhance policy learning in both robot manipulation and navigation tasks. (iii) We conduct extensive experiments, especially with other possible termination conditions, to provide insights for future research in the field of skill termination learning.
Figure 1: Visualization of an example of critical decision points in the kitchen environment. High state-action novelty can be found in states where a subtask has been completed, and multiple subsequent subtasks are accessible.
@@ 2 Related Works
One major approach of discovering good options is to focus on identifying good terminal states, or sub-goal states. For example, landmark states (Kaelbling, 1993), reinforcement learning signals (Digney, 1998), graph partitioning (Menache et al., 2002; Şimşek et al., 2005; Machado et al., 2017a; b), and state clustering (Srinivas et al., 2016) have been used to identify meaningful sub-goal states. Digney (1998); Simsek et al. (2005) and Kulkarni et al. (2016) focused on detecting bottleneck states, which are states that appear frequently within successful trajectories, but are less common in unsuccessful trajectories (e.g., a state with access door). Şimşek & Barto (2004) tried to identify access states, which are similar to bottleneck states, but determined based on the relative novelty score of predecessor states and successor states. Access states are found based on the intuition that sub-goals will exhibit a relative novelty score distribution with scores that are frequently higher than those of non sub-goals. These studies motivated us to search for states with meaningful properties to terminate skills. However, these methods frequently face challenges in scaling to large or continuous state spaces.
@@
As extending the classic option framework to high-dimensional state spaces through the adoption of function approximation is not straightforward, a number of practitioners have proposed acquiring skills, which are fixed-length executions of low-level policies, to achieve temporal abstraction. For example, skill discovery (Gregor et al., 2016; Achiam et al., 2018; Mavor-Parker et al., 2022; Park et al., 2023) and skill extraction (Yang et al., 2021; Singh et al., 2020; Pertsch et al., 2021b; Hakhamaneshi et al., 2021) frameworks have proven to be successful in acquiring meaningful sets of skills. Especially, Pertsch et al. (2021a) showed promising results in complex, long-horizon tasks with sparse rewards by extracting skills with data-driven behavior priors. The learned prior enables the agent to explore the environment in a more structured manner, which leads to better performance in downstream tasks. However, we believe that their performances are greatly constrained by the use of fixed-length skills, which restricts them from making decisions at critical decision points. There are prior works introducing variable-length skill extraction methods. Salter et al. (2022) focus on learning an option-level transition model leveraging predictability to compress offline behaviors into options that terminate at bottleneck states. However, due to its model-based design, the extracted skills are not suitable for transferring to downstream tasks with significantly different environment configuration. Jiang et al. (2022) is an option framework based model that learns both options and termination condition from the task-agnostic demonstrations in terms of minimum description length (Rissanen, 1978). As it serves as an appropriate benchmark for evaluating our approach, we selected Jiang et al. (2022) as one of the baselines to compare with.
@@ 4 Simple and Effective Identification of Decision Points
The option framework aims to achieve temporal abstraction by learning good options, and good options can be learned through the identification of meaningful sub-goal states (Menache et al., 2002; Şimşek & Barto, 2004), i.e., the critical decision points. In this work, we propose to use state-action novelty to identify critical decision points for skill termination, which leads to the execution of variable-length skills. In particular, we use intrinsic curiosity module (ICM) (Pathak et al., 2017) as our state-action novelty estimator (more details in Section G.2). However, any state-action novelty estimation mechanism that measures the joint novelty of state-action pairs can be used for our approach.
@@ 4.1 State-action Novelty-based Decision Point Identification
Figure 2: Visualization of prediction error of ICM in maze and block stacking environment. Note the same offline data that is used to train ICM was used to compute this prediction error. (A), (B) and (C) are the state-action pairs with the highest prediction error, while (D), (E) and (F) are the ones with the lowest.
@@
In the kitchen environment, as shown in Figure 1, high state-action novelty $\chi(s,a)$ tends to occur in states where a subtask has been completed. After completing the subtask of flipping the light switch, the agent has the option to open either the left hinge cabinet or the right slide cabinet. In sequential manipulation tasks, such critical points are valuable because completing one subtask grants access to multiple other subtasks.
Figure 3: The relative frequency of termination improvement occurrences (left), conditional action novelty (middle), and state novelty (right) in a small grid with three different goals. Higher percentile colors indicate a relatively greater number of termination improvement occurrences, higher conditional action novelty, and higher state novelty. Further details on the visualization procedure are provided in Appendix G.1.
@@ 4.2 Termination Improvement from State-action Novelty-based Terminations
[Termination Improvement, Sutton (1998), informal] For any meta-control policy $\mu$ on set of options $\mathcal{O}$, define a new set of options $\mathcal{O^{\prime}}$, which is a set of options that we can additionally choose to terminate whenever the value of a state $V^{\mu}(s)$ is larger than the value of a state given that we keep the current option $o$, $Q^{\mu}(s,o)$. With $\mu^{\prime}$, which has the same option selection probability as $\mu$ but over a new set of options $\mathcal{O^{\prime}}$, we have $V^{\mu^{\prime}}(s)\geq V^{\mu}(s)$.
The termination improvement theorem implies that we should terminate an option when there are much better alternatives available from the current state. When the skills (or options) are discovered from diverse trajectories (e.g., trajectories gathered from a diverse set of goals), termination improvement is typically observed in states where a multitude of actions have been executed, such as crossroads.
To identify the states where termination improvement occurs, we plotted the relative frequency of termination improvement occurrences in a small $8\times 8$ grid maze with three different goal settings (Figure 3 (left)). It shows that termination improvement frequently occurs in states where diverse plausible actions exist. In states with a single available option, $V^{\mu}(s)$ would be equal to $Q^{\mu}(s,o)$. On the other hand, as more actions/options are plausible, $Q^{\mu}(s,o)$ would exhibit a broader range of values, thereby increasing the likelihood of satisfying $Q^{\mu}(s,o)<V^{\mu}(s)$.
@@ 5 Learning Termination Conditions through State-action Novelty Module
In this work, we propose to use state-action novelty as a termination condition to effectively capture critical decision points and execute terminated skills. Our approach consists of two major steps. First, we train the state-action novelty module and then the low-level policy using an offline trajectory dataset for skill extraction. Next, we perform online reinforcement learning with the pre-trained state-action novelty module to solve an unseen task.
@@
For training the state-action novelty module, we assume access to unstructured agent experiences of states and actions in the form of $N$ trajectories $\mathcal{D}=\left\{\tau^{i}=\{(s_{t},a_{t})\}_{t=0}^{T-1}\right\}_{i=0}^{N-1}$. These trajectories are collected across a diverse set of tasks except for the one we are specifically interested in (see Appendix E for more details). Since we do not make any assumptions about rewards or task labels, our model can leverage real-world datasets that can be collected at a lower cost (e.g., autonomous driving and drones).
@@ 5.1 Unsupervised Learning of State-action Novelty Module
In the process of unsupervised learning, our goal is to pre-train the low-level policy $\pi(a|z,s)$ and the state-action novelty module $p(\beta|s,a)$. We define a skill $z\in\mathcal{Z}$ as an embedding of state-action pairs $\tau=\{(s_{i},a_{i})\}_{i=t}^{t+H-1}$ and termination conditions $\boldsymbol{\beta}=\{\beta_{i}\}_{i=t}^{t+H-1}$. The termination conditions $\beta$ are Bernoulli random variables that decide when to stop the current skill. Through the classification of state-action pairs demonstrating significant novelty $\chi(s,a)$, $\beta$ are trained to predict the critical decision points. The point at which novelty is considered significant varies depending on the environment. In downstream tasks, the skill being executed will be terminated either when $\beta=1$ is sampled or when the maximum skill length $H$ is reached. In all experiments, we set $H=30$. The low-level policy is trained following standard practices (Pertsch et al., 2021a; Hakhamaneshi et al., 2021) (see Appendix B for details).
5.2 Reinforcement Learning with Skill Termination Conditions
@@ 6.1 State-action Novelty Module
6.2 Environments
A large set of task-agnostic agent experiences is collected from each environment to pre-train the low-level policy and the state-action novelty module. For downstream tasks with significantly different environment configurations (maze, sparse block stacking), we used agent-centered cropped images to extract consistent structural information from task-agnostic demonstrations (see Appendix E). We evaluate the models on their performance in solving complex, unseen tasks in each environment. Further details of the environments and the task-agnostic data collection procedure are in Appendix F.
6.3 Results
We use the following models for comparison: Flat RL (SAC): Soft Actor-Critic (Haarnoja et al., 2018) agent that does not leverage prior experience. This comparison illustrates the effectiveness of temporal abstraction. SAC+Novelty: Flat RL that uses state-action novelty as intrinsic rewards. This comparison demonstrates the importance of incorporating state-action novelty in skill learning. Flat Offline Learning w/ Finetuning (BC+SAC, IQL+SAC): The supervised behavioral cloning (BC) policy and Implicit Q-Learning (IQL) (Kostrikov et al., 2021) policy that are trained on offline data and subsequently fine-tuned for the downstream task using SAC. Fixed-length Skill Policy (SPiRL): The agent that learns a fixed-length skill policy (Pertsch et al., 2021a) by leveraging prior experience. This comparison demonstrates the benefit of critical decision points identification through state-action novelty. NBDI (Ours): The agent that learns a terminated skill policy through state-action novelty $\chi(s,a)$. It learns a state-action novelty based termination distribution $p(\beta|z,s)$ to predict skill termination at current step. State Novelty Decision Point Identification (NBDI-$\chi(s)$): The agent that learns a terminated skill policy through state novelty. To exclusively assess the influence of the novelty type, we distilled the state-action novelty module used in NBDI into a separate network, $\chi(s)$, which solely depends on the current state. Conditional Action Novelty Decision Point Identification (NBDI-$\chi(a|s)$): The agent that learns a terminated skill policy through conditional action novelty $\frac{\chi(s,a)}{\chi(s)}$, where $\chi(s)$ is the distilled state novelty module used for NBDI-$\chi(s)$. Implementation details of the baselines are in Appendix G.
@@ 6.3 Results
6.4 Comparison to Other Variable-length Skill Extraction Methods
We compare the performance of NBDI, relative novelty (Şimşek & Barto, 2004) and LOVE (Jiang et al., 2022) in the maze and kitchen environment (Figure 6(b)). Relative novelty identifies termination condition based on the assumption that sub-goals typically show a relative novelty score distribution with higher scores than those of non-sub-goals. LOVE is an option framework based model that learns both options and termination condition from the task-agnostic demonstrations in terms of minimum description length. Since LOVE is designed for discrete action space, we used an MLP layer as the action decoder to handle continuous actions. Furthermore, since LOVE was only evaluated in downstream tasks that has the same MDP as the data collection environment (except the reward function), we evaluated whether it can transfer to downstream tasks with significantly different environment configuration (Maze $40\times 40$). Implementation details of baselines are in Appendix G.4.

### v3

@@ 1 Introduction
In this paper, we present NBDI (Novelty-based Decision Point Identification) [^1] Code:  https://github.com/ku-dmlab/NBDI, a simple state-action novelty-based decision point identification method that allows the agent to learn terminated skills from task-agnostic demonstrations. In this context, the term task-agnostic refers to the collection of trajectories from a diverse set of tasks, excluding the one we are specifically interested in (see Appendix E for visualizations). Identifying critical decision points promote knowledge transfer between different tasks and stimulate exploration by closely connecting different areas in the state space (McGovern & Barto, 2001; Menache et al., 2002; Şimşek & Barto, 2004). For example, detecting doorways between rooms is useful regardless of the specific task at hand. We demonstrate the straightforward applicability of our method to the skill-based deep RL framework, and illustrate how it can lead to improvements in decision-making. To summarize, our three main contributions are as follows: (i) For the first time, we propose a skill termination condition for task-agnostic demonstrations that remains effective even when the environment configuration of complex, long-horizon downstream tasks undergo significant changes. (ii) We present a novel method in reinforcement learning, which is the identification of state-action novelty-based critical decision points. Furthermore, we demonstrate that executing terminated skills, based on state-action novelty, can substantially enhance policy learning in both robot manipulation and navigation tasks. (iii) We conduct extensive experiments, especially with other possible termination conditions, to provide insights for future research in the field of skill termination learning.
Figure 1: Visualization of an example of critical decision points in the kitchen environment. High state-action novelty can be found in states where a subtask has been completed, and multiple subsequent subtasks are accessible. If termination occurs at a high state-action novelty point, the agent retains multiple plausible options. For example, after moving the kettle, it may choose to open the sliding cabinet or turn the oven knob (Left). Similarly, after flipping the light switch, the agent may proceed to open either the sliding cabinet or the left-hinged cabinet (Right).
@@ 2 Related Works
One major approach of discovering good options is to focus on identifying good terminal states, or sub-goal states. For example, landmark states (Kaelbling, 1993), reinforcement learning signals (Digney, 1998), graph partitioning (Menache et al., 2002; Şimşek et al., 2005; Machado et al., 2017a, b), and state clustering (Srinivas et al., 2016) have been used to identify meaningful sub-goal states. (Digney, 1998; Simsek et al., 2005) and (Kulkarni et al., 2016) focused on detecting bottleneck states, which are states that appear frequently within successful trajectories, but are less common in unsuccessful trajectories (e.g., a state with access door). (Şimşek & Barto, 2004) tried to identify access states, which are similar to bottleneck states, but determined based on the relative novelty score of predecessor states and successor states. Access states are found based on the intuition that sub-goals will exhibit a relative novelty score distribution with scores that are frequently higher than those of non sub-goals. These studies motivated us to search for states with meaningful properties to terminate skills. However, these methods frequently face challenges in scaling to large or continuous state spaces.
@@
As extending the classic option framework to high-dimensional state spaces through the adoption of function approximation is not straightforward, a number of practitioners have proposed acquiring skills, which are fixed-length executions of low-level policies, to achieve temporal abstraction. For example, skill discovery (Gregor et al., 2016; Achiam et al., 2018; Mavor-Parker et al., 2022; Park et al., 2023) and skill extraction (Yang et al., 2021; Singh et al., 2020; Pertsch et al., 2021b; Hakhamaneshi et al., 2021) frameworks have proven to be successful in acquiring meaningful sets of skills. Especially, Pertsch et al. (2021a) showed promising results in complex, long-horizon tasks with sparse rewards by extracting skills with data-driven behavior priors. The learned prior enables the agent to explore the environment in a more structured manner, which leads to better performance in downstream tasks. However, we believe that their performances are greatly constrained by the use of fixed-length skills, which restricts them from making decisions at critical decision points. There are prior works introducing variable-length skill extraction methods. Salter et al. (2022) focus on learning an option-level transition model leveraging predictability to compress offline behaviors into options that terminate at bottleneck states. However, due to its model-based design, the extracted skills are not suitable for transferring to downstream tasks with significantly different environment configuration. Jiang et al. (2022) is an option framework based model that learns both options and termination condition from the task-agnostic demonstrations in terms of minimum description length (Rissanen, 1978). As it serves as a suitable benchmark for evaluating our approach, we later compare our method to Jiang et al. (2022) in Section 6.4.
@@ 4 Simple and Effective Identification of Decision Points
The option framework aims to achieve temporal abstraction by learning good options, and good options can be learned through the identification of meaningful sub-goal states (Menache et al., 2002; Şimşek & Barto, 2004), i.e., the critical decision points. In this work, we propose to use state-action novelty to identify critical decision points for skill termination, which leads to the execution of variable-length skills. In particular, we use intrinsic curiosity module (ICM) (Pathak et al., 2017) as our state-action novelty estimator (more details in Section I.2). However, any state-action novelty estimation mechanism that measures the joint novelty of state-action pairs can be used for our approach.
@@ 4.1 State-action Novelty-based Decision Point Identification
Figure 2: Visualization of prediction error of ICM in maze and block stacking environment. Note the same offline data that is used to train ICM was used to compute this prediction error. (A), (B) and (C) are the state-action pairs with the highest prediction error, while (D), (E) and (F) are the ones with the lowest. Critical decision points—such as crossroads or states involving block manipulation—are typically associated with high prediction error. In contrast, low prediction errors are observed in states that are less important for decision-making.
@@
In the kitchen environment, as shown in Figure 1, high state-action novelty $\chi(s,a)$ tends to occur in states where a subtask has been completed. For example, after moving the kettle, the agent may choose to either open the sliding cabinet or turn the oven knob. Similarly, after completing the subtask of flipping the light switch, the agent has the option to open either the left hinge cabinet or the right slide cabinet.
In sequential manipulation tasks, such critical points are valuable because completing one subtask grants access to multiple other subtasks.
Figure 3: The relative frequency of termination improvement occurrences (left), conditional action novelty (middle), and state novelty (right) in a small grid with three different goals. Higher percentile colors indicate a relatively greater number of termination improvement occurrences, higher conditional action novelty, and higher state novelty. Further details on the visualization procedure are provided in Appendix I.1.
@@ 4.2 Termination Improvement from State-action Novelty-based Terminations
[Termination Improvement, (Sutton, 1998), informal] For any meta-control policy $\mu$ on set of options $\mathcal{O}$, define a new set of options $\mathcal{O^{\prime}}$, which is a set of options that we can additionally choose to terminate whenever the value of a state $V^{\mu}(s)$ is larger than the value of a state given that we keep the current option $o$, $Q^{\mu}(s,o)$. With $\mu^{\prime}$, which has the same option selection probability as $\mu$ but over a new set of options $\mathcal{O^{\prime}}$, we have $V^{\mu^{\prime}}(s)\geq V^{\mu}(s)$.
The termination improvement theorem implies that we should terminate an option when there are much better alternatives available from the current state.
To identify the states where termination improvement occurs, we plotted the relative frequency of termination improvement occurrences in a small $8\times 8$ grid maze with three different goal settings (Figure 3 (left)). It shows that termination improvement frequently occurs in states where diverse plausible actions exist. In states with a single available option, $V^{\mu}(s)$ would be equal to $Q^{\mu}(s,o)$. On the other hand, as more actions/options are plausible, $Q^{\mu}(s,o)$ would exhibit a broader range of values, thereby increasing the likelihood of satisfying $Q^{\mu}(s,o)<V^{\mu}(s)$. When the skills (or options) are discovered from diverse trajectories (e.g., trajectories gathered from a diverse set of goals), termination improvement is typically observed in states where a multitude of actions have been executed, such as crossroads.
@@ 5 Learning Termination Conditions through State-action Novelty Module
In this work, we propose to use state-action novelty as a termination condition to effectively capture critical decision points and execute terminated skills. Our approach consists of two major steps. First, we train the state-action novelty module and then the low-level policy using task-agnostic demonstrations for skill extraction. Next, we perform online reinforcement learning with the learned variable-length skills to solve an unseen task.
@@
For training the state-action novelty module, we assume access to task-agnostic, expert-level demonstrations of states and actions in the form of $N$ trajectories $\mathcal{D}=\left\{\tau^{i}=\{(s_{t},a_{t})\}_{t=0}^{T-1}\right\}_{i=0}^{N-1}$. These trajectories are collected across a diverse set of tasks except for the one we are specifically interested in (see Appendix E for more details). Since we do not make any assumptions about rewards or task labels, our model can leverage real-world datasets that can be collected at a lower cost (e.g., autonomous driving and drones).
@@ 5.1 Unsupervised Learning of State-action Novelty Module
In the process of unsupervised learning, our goal is to pre-train the low-level policy $\pi(a|z,s)$ and the state-action novelty module $p(\beta|s,a)$. We define a skill $z\in\mathcal{Z}$ as an embedding of state-action pairs $\tau=\{(s_{i},a_{i})\}_{i=t}^{t+H-1}$ and termination conditions $\boldsymbol{\beta}=\{\beta_{i}\}_{i=t}^{t+H-1}$. The termination conditions $\beta$ are Bernoulli random variables that decide when to stop the current skill. Through the classification of state-action pairs demonstrating significant novelty $\chi(s,a)$, $\beta$ are trained to predict the critical decision points. The point at which novelty is considered significant varies depending on the environment. In downstream tasks, the skill being executed will be terminated either when $\beta=1$ is sampled or when the maximum skill length $H$ is reached. In all experiments, we set $H=30$. The low-level policy is trained following standard practices (Pertsch et al., 2021a; Hakhamaneshi et al., 2021). During the training of the low-level policy, the deep latent variable model receives a randomly sampled experience from the training dataset, along with a termination condition vector provided by the state-action novelty module. The model is trained to reconstruct the corresponding action sequence and its length (i.e., the termination point) by maximizing the evidence lower bound (ELBO) (see Appendix B for details).
@@
5.2 Reinforcement Learning with Skill Termination Conditions
@@ 6.1 State-action Novelty Module
6.2 Environments
A large set of task-agnostic agent experiences is collected from each environment to pre-train the low-level policy and the state-action novelty module. For downstream tasks with significantly different environment configurations (maze, sparse block stacking), we used agent-centered cropped images to extract consistent structural information from task-agnostic demonstrations. We introduce the train/transfer domain similarity for the environments used in our experiments in Appendix E. To demonstrate the effectiveness of our method on complex downstream tasks involving significant configuration changes, we evaluate the models on a maze environment with an unseen layout and a large-scale block stacking environment with a greater number of blocks in randomized positions. Both settings differ significantly from the task-agnostic offline datasets. Further details of the environments and the task-agnostic data collection procedure are in Appendix H.
6.3 Results
We use the following models for comparison: Flat RL (SAC): Soft Actor-Critic (Haarnoja et al., 2018) agent that does not leverage prior experience. This comparison illustrates the effectiveness of temporal abstraction. SAC+Novelty: Flat RL that uses state-action novelty as intrinsic rewards. This comparison demonstrates the importance of incorporating state-action novelty in skill learning. Flat Offline Learning w/ Finetuning (BC+SAC, IQL+SAC): The supervised behavioral cloning (BC) policy and Implicit Q-Learning (IQL) (Kostrikov et al., 2021) policy that are trained on offline data and subsequently fine-tuned for the downstream task using SAC. Fixed-length Skill Policy (SPiRL): The agent that learns a fixed-length skill policy (Pertsch et al., 2021a) by leveraging prior experience. This comparison demonstrates the benefit of critical decision points identification through state-action novelty. NBDI (Ours): The agent that learns a terminated skill policy through state-action novelty $\chi(s,a)$. It learns a state-action novelty based termination distribution $p(\beta|z,s)$ to predict skill termination at current step. State Novelty Decision Point Identification (NBDI-$\chi(s)$): The agent that learns a terminated skill policy through state novelty. To exclusively assess the influence of the novelty type, we distilled the state-action novelty module used in NBDI into a separate network, $\chi(s)$, which solely depends on the current state. Conditional Action Novelty Decision Point Identification (NBDI-$\chi(a|s)$): The agent that learns a terminated skill policy through conditional action novelty $\frac{\chi(s,a)}{\chi(s)}$, where $\chi(s)$ is the distilled state novelty module used for NBDI-$\chi(s)$. Implementation details of the baselines are in Appendix I.
@@ 6.3 Results
6.4 Comparison to Other Variable-length Skill Extraction Methods
We compare the performance of NBDI, relative novelty (Şimşek & Barto, 2004) and LOVE (Jiang et al., 2022) in the maze and kitchen environment (Figure 6(b)). Relative novelty identifies termination condition based on the assumption that sub-goals typically show a relative novelty score distribution with higher scores than those of non-sub-goals. LOVE is an option framework based model that learns both options and termination condition from the task-agnostic demonstrations in terms of minimum description length. Since LOVE is designed for discrete action space, we used an MLP layer as the action decoder to handle continuous actions. Furthermore, since LOVE was only evaluated in downstream tasks that has the same MDP as the data collection environment (except the reward function), we evaluated whether it can transfer to downstream tasks with significantly different environment configuration (Maze $40\times 40$). Implementation details of baselines are in Appendix I.4.
Figure 8: Illustration of the impact of varying the width and depth configuration of hidden layer and dataset usage on state-action novelty estimation. (Left) shows how varying hidden layer configurations (depth and width) influences the estimated state-action novelty (box plots) and the total number of parameters (blue line). The x-axis lists different hidden layer setups. (Right) illustrates how varying the proportion of dataset usage affects novelty estimation.
Figure 9: Comparison of success rates in the maze environment during the skill-based meta-training phase of SiMPL across 10 meta-training tasks (10 random goal locations). The shaded regions indicate standard errors over three random seeds.